OpenStudy (anonymous):

Sarah needs to make a cake and some cookies. The cake requires 3/8 cup of sugar and the cookies require 3/5 cup of sugar. Sarah has 15/16 cups of sugar. Does she have enough sugar, or how much more does she need?

OpenStudy (anonymous):

So in order to do this problem you will need to make the denominators like terms because this is an addition problem.

OpenStudy (anonymous):

30?

OpenStudy (anonymous):

3/8 + 3/5 = 15/16 ... would be the equation to this

OpenStudy (anonymous):

The common denominator cannot be 30 because 16 and 8 don't multiply evenly into it.

OpenStudy (anonymous):

Let's take this in chunks. Can you think of anything common that 5 and 8 can multiply into?

OpenStudy (anonymous):

40 ...

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i was getting 6 x5 and 8x5 backwards

OpenStudy (anonymous):

Haha, no worries. Do you know how to take 3/8 and turn it into a fraction over 40?

OpenStudy (anonymous):

yeah 8 goes into 40 5times so 5 times 3 =15 15/40

OpenStudy (anonymous):

Exactly. Now do the same for the 3/5?

OpenStudy (anonymous):

24/40

OpenStudy (anonymous):

Perfect. So now we can actually add. 15/40 + 24/40 = ?

OpenStudy (anonymous):

actually while we are at it ... how do you figure out the smallest common denominator? i always had a hard time with that. wasnt it something about prime numbers?

OpenStudy (anonymous):

or would it be better to make it a new question?

OpenStudy (anonymous):

39/40

OpenStudy (anonymous):

Oh no, I can answer it. If it takes you a while to figure it out, it's more than likely the lowest common denominator is some random number we don't know off the tops of our heads. I wouldn't bother with it in that case, just use 8*5*16 as the common denominator and then work from there.

OpenStudy (anonymous):

ok so how much more sugar does she need and how do we determine that?

OpenStudy (anonymous):

subract 39/40 and 15/16

OpenStudy (anonymous):

wasnt there a simple rule to figure it out? i remember me doing this in 6th grade i think.

OpenStudy (anonymous):

You have to figure out if 39/40 < 15/16, and if it is, she doesn't have enough.

OpenStudy (anonymous):

So you must compare with like denominators. Once you have a common denominator for 40 and 16, you can compare with ease.

OpenStudy (anonymous):

I'm really sorry, but I'm at school and the period is up. I have to go.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

And yes, she does have enough ;)

OpenStudy (anonymous):

.975 > .9375

OpenStudy (anonymous):

See you later!